# Class Distribution Shifts in Zero-Shot Learning: Learning Robust Representations

### v3

@@ 6 Discussion
In this study, we examined class distribution shifts in zero-shot learning, with a focus on shifts induced by unknown attributes. Such shifts pose significant challenges in zero-shot learning, where new classes emerge in testing, causing standard techniques trained via ERM to fail on shifted class distributions, even when the conditional distribution of the data given the class remains the same.
Existing lines of work (see Appendix A) assumes closed-world classification or a known cause, making it unsuitable for zero-shot learning or shifts caused by unknown attributes. In response, we introduced a straightforward framework and the first algorithm to address class distribution shifts in zero-shot learning using OOD environment balancing methods.
In the causal terminology of closed-world OOD generalization, our framework employs synthetic environments to intervene on attribute mixtures by sampling small class subsets, thereby manipulating the class distribution. This facilitates the creation of diverse environments with varied attribute mixtures, enhancing the distinction between negative examples. We further discuss the comparison with OOD balancing methods in Appendix C. Additionally, our proposed VarAUC penalty, designed for metric losses, enhances the separation of negative examples.
Our results demonstrate improvements compared to the ERM baseline on shifted distributions without compromising performance on unshifted ones, enabling the learning of more robust representations for zero-shot tasks and ensuring reliable performance.
While the proposed framework is general, our current experiments address shifts in a binary attribute. We defer exploration of additional scenarios, such as those involving shifts in multiple correlated attributes, to future work. An additional promising direction for future work is the consideration of shifts where the responsible attribute is strongly correlated with additional attributes or covariates. This opens possibilities to explore structured constructions of synthetic environments that leverages such correlations.
@@ Appendix A Related Work on Class Distribution Shifts in Closed-World Settings
In Class-Imbalanced Learning [28, 10, 8] it is assumed that some classes are more dominant in training, while in deployment this is no longer the case. Therefore, solutions classically include data or loss re-weighting [46, 6, 40, 32] and calibration of the classification score [44, 60]. A particularly popular framework to address class distribution shifts is Distributionally Robust Optimization (DRO) [3, 13, 12, 55], where instead of assuming a specific probability distribution, a set or a range of possible distributions is considered, and optimization is performed to achieve best results over the worst-case distribution. A special case known as Group DRO [45, 38], involves a group variable that introduces discriminatory patterns among classes within specific groups. This framework includes methods that assume that the classifier does not have access to the group information, and therefore propose re-weighting high loss examples [29], and data sub-sampling to balance classes and groups [21]. Nevertheless, all these methodologies rely on the training and test class sets being identical, making them unsuitable for direct application in zero-shot learning scenarios.
@@ B.1 Derivation of the Loss
We begin by revisiting the parametric model introduced in §3. Let $z_{i}|c_{i}\sim\mathcal{N}\left(c_{i},\Sigma_{z}\right)$, where $\Sigma_{z}=\nu_{z}I_{d}$, $0<\nu_{z}\in\mathbb{R}$ and $I_{d}$ is the $d$ dimensional identity matrix. Classes $c_{i}$ are drawn according to a Gaussian distribution $c_{i}\sim\mathcal{N}\left(0,\Sigma_{a}\right)$ corresponding to their type $a\in\{a_{1},a_{2}\}$. Here, we use a simpler (although less intuitive) notation for the values of the diagonal matrices $\Sigma_{a}$:
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left\|w\left(z_{i}-z_{j}\right)\right\|^{4}$ $=\operatorname{\mathbb{E}}_{y_{ij}=0}\left[2\operatorname{tr}\left(\Sigma^{2}\right)+4\operatorname{tr}\left(\Sigma\Sigma_{\mu}\right)+\operatorname{tr}^{2}(\Sigma)+2\operatorname{tr}\left(\Sigma\right)\operatorname{tr}\left(\Sigma_{\mu}\right)+2\operatorname{tr}\left(\Sigma_{\mu}^{2}\right)+\left(\operatorname{tr}\left(\Sigma_{\mu}\right)\right)^{2}\right]$
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}\left(\Sigma\Sigma_{\mu}\right)\right]$ $=2\nu_{z}\operatorname{tr}\left(W^{T}W\left[2\rho^{2}W^{T}\Sigma_{a_{1}}W+2(1-\rho)^{2}W^{T}\Sigma_{a_{2}}W+2\rho(1-\rho)W^{T}\left(\Sigma_{a_{1}}+\Sigma_{a_{2}}\right)W\right]\right)$
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}\left(\Sigma_{\mu}\right)\right]$ $=\rho^{2}\operatorname{tr}\left(2W^{T}\Sigma_{a_{1}}W\right)+(1-\rho)^{2}\operatorname{tr}\left(2W^{T}\Sigma_{a_{1}}W\right)+2\rho(1-\rho)\operatorname{tr}\left(W^{T}\left(\Sigma_{a_{1}}+\Sigma_{a_{2}}\right)W\right)$
$=2\left[\nu_{0}\sum_{i=1}^{d_{0}}w_{i}^{2}+\alpha_{1}\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}+\alpha_{2}\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}\right],$ (27)
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}\left(\Sigma_{\mu}^{2}\right)\right]$ $=\rho^{2}\operatorname{tr}\left((2W^{T}\Sigma_{a_{1}}W)^{2}\right)+(1-\rho)^{2}\operatorname{tr}\left((2W^{T}\Sigma_{a_{2}}W)^{2}\right)+2\rho(1-\rho)\operatorname{tr}\left((W^{T}\left(\Sigma_{a_{1}}+\Sigma_{a_{2}}\right)W)^{2}\right)$
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}^{2}\left(\Sigma_{\mu}\right)\right]$ $=2\left[2\nu_{0}^{2}\left(\sum_{i=1}^{d_{0}}w_{i}^{2}\right)^{2}+\beta_{1}\left(\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}\right)^{2}+\beta_{2}\left(\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}\right)^{2}\right.$
$+4\gamma_{0,1}\sum_{i=1}^{d_{0}}w_{i}^{2}\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}+4\gamma_{0,2}\sum_{i=1}^{d_{0}}w_{i}^{2}\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}+4\gamma_{1,2}\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}\right],$ (30)
@@
Finally, to symmetry, at the optimal solution we have
$w_{i}=\begin{cases}u_{0}&0\leq i\leq d_{0}\\ u_{1}&d_{0}+1\leq i\leq d_{0}+d_{1}\\ u_{2}&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (32)
@@ B.2 Analysis of the Optimal Solution (Proof of Proposition 1)
Proposition 1 shows that for not too large values of $d_{2}$, the optimal solution on the training distribution, assigns higher weight to components corresponding to high variance for training data, than to components corresponding to high variance on the shifted test distribution.
We begin by defining the required condition on $d_{2}$. Denote
@@
$h\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2},d_{1}\right)\coloneqq\max\left\{(2+d_{1})\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}}-2,(2+d_{1})\frac{\psi_{1}}{\psi_{2}}\frac{\alpha_{2}+\nu_{z}}{\alpha_{1}+\nu_{z}}-2,\frac{\eta_{12}}{2\psi_{2}}\right\}.$ (35)
The condition on $d_{2}$ in Proposition 1 is that $d_{2}\leq h\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2},d_{1}\right)$. This provides a sufficient but not necessary condition – this is a simple example that does not require a full characterization of the optimal solution, which is provided in Appendix B.3.
@@
Since $\eta_{02}\leq\eta_{01}$ the assumption $\left(2+d_{2}\right)\psi_{2}\eta_{01}\leq\left(2+d_{1}\right)\psi_{1}\eta_{02}$ also yields $\left(2+d_{2}\right)\psi_{2}\leq\left(2+d_{1}\right)\psi_{1}$. Hence, for
$d_{2}+2<\max\left\{(2+d_{1})\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}},(2+d_{1})\frac{\psi_{1}}{\psi_{2}}\frac{\alpha_{2}+\nu_{z}}{\alpha_{1}+\nu_{z}},\frac{\eta_{12}}{2\psi_{2}}\right\}$
all the terms in B.2 are positive and therefore $d_{1}u_{1}^{2}-d_{2}u_{2}^{2}>0$. ∎
@@ B.3 Explicit Expression for the Optimal Representation
$\frac{\partial}{\partial\left(w_{i}^{2}\right)}\operatorname{\mathbb{E}}_{y=0}\left[\operatorname{tr}\left(\Sigma\Sigma_{\mu}\right)\right]$ =$=\begin{cases}8\nu_{z}\nu_{0}w_{i}^{2}&1\leq i\leq d_{0}\\ 8\nu_{z}\alpha_{1}w_{i}^{2}&d_{0}+1\leq i\leq d_{0}+d_{1}\\ 8\nu_{z}\alpha_{2}w_{i}^{2}&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (45)
$\frac{\partial\big{[}\operatorname{tr}(\Sigma)\operatorname{\mathbb{E}}_{y=0}[\operatorname{tr}\Sigma_{\mu}]\big{]}}{\partial\left(w_{i}^{2}\right)}$ =$=\begin{cases}4\nu_{z}\left[2\nu_{0}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+\left(\alpha_{1}+\nu_{0}\right)\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+\left(\alpha_{2}+\nu_{0}\right)\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}\right]&1\leq i\leq d_{0}\\ 4\nu_{z}\left[\left(\nu_{0}+\alpha_{1}\right)\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+2\alpha_{1}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+\left(\alpha_{2}+\alpha_{1}\right)\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}\right]&d_{0}+1\leq i\leq d_{0}+d_{1}\\ 4\nu_{z}\left[\left(\nu_{0}+\alpha_{2}\right)\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+\left(\alpha_{1}+\alpha_{2}\right)\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+2\alpha_{2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}\right]&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (46)
@@
$\frac{\partial}{\partial\left(w_{i}^{2}\right)}\operatorname{\mathbb{E}}_{y=0}\left[\operatorname{tr}^{2}\left(\Sigma_{\mu}\right)\right]$ =$=\begin{cases}8\nu_{0}^{2}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+8\gamma_{0,1}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+8\gamma_{0,2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}&1\leq i\leq d_{0}\\ 8\gamma_{0,1}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+4\beta_{1}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+8\gamma_{1,2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}&d_{0}+1\leq i\leq d_{0}+d_{1}\\ 8\gamma_{0,2}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+8\gamma_{1,2}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+4\beta_{2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (48)
@@
and thus we can write for the symmetric solution
$\partial_{0}$ $=-2m\left(\nu_{0}+\nu_{z}\right)+u_{0}^{2}A_{0,0}+u_{1}^{2}A_{0,1}+u_{2}^{2}A_{0,2},$ (49)
$\partial_{1}$ $=-2m\left(\alpha_{1}+\nu_{z}\right)+u_{0}^{2}A_{1,0}+u_{1}^{2}A_{1,1}+u_{2}^{2}A_{1,2},$ (50)
$\partial_{2}$ $=-2m\left(\alpha_{2}+\nu_{z}\right)+u_{0}^{2}A_{2,0}+u_{1}^{2}A_{2,1}+u_{2}^{2}A_{2,2},$ (51)

### v4

@@ 6 Discussion
In this study, we examined class distribution shifts in zero-shot learning, with a focus on shifts induced by unknown attributes. Such shifts pose significant challenges in zero-shot learning where new classes emerge in testing, causing standard techniques trained via ERM to fail on shifted class distributions, even when the conditional distribution of the data given class remains the same.
Previous research (see Appendix A) assumes closed-world classification or a known cause, making these methods unsuitable for zero-shot learning or shifts caused by unknown attributes. In response, we introduced a framework and the first algorithm to address class distribution shifts in zero-shot learning using OOD environment balancing methods.
In the causal terminology of closed-world OOD generalization, our framework employs synthetic environments to intervene on attribute mixtures by sampling small class subsets, thereby manipulating the class distribution. This facilitates the creation of diverse environments with varied attribute mixtures, enhancing the distinction between negative examples. A further comparison of our framework with OOD environment balancing methods is provided in Appendix C. Additionally, our proposed VarAUC penalty, designed for metric losses, enhances the separation of negative examples.
Our results demonstrate improvements compared to the ERM baseline on shifted distributions, without compromising performance on unshifted distributions, enabling the learning of more robust representations for zero-shot tasks and ensuring reliable performance.
While the proposed framework is general, our current experiments address shifts in a binary attribute. We defer exploration of additional scenarios, such as those involving shifts in multiple correlated attributes, to future work. An additional promising direction for future work is the consideration of shifts where the responsible attribute is strongly correlated with additional attributes or covariates. This opens up possibilities to explore structured constructions of synthetic environments that leverage such correlations.
@@ Appendix A Related Work on Class Distribution Shifts in Closed-World Settings
In class-imbalanced learning [28, 10, 8] it is assumed that some classes are more dominant in training, while in deployment this is no longer the case. Therefore, solutions classically include data or loss re-weighting [46, 6, 40, 32] and calibration of the classification score [44, 60]. A popular framework for addressing class distribut ion shifts is distributionally robust optimization (DRO) [3, 13, 12, 55], where instead of assuming a specific probability distribution, a set or range of possible distributions is considered, and optimization is performed to achieve the best results on the worst-case distribution. A special case known as group DRO [45, 38], involves a group variable that introduces discriminatory patterns among classes within specific groups. The framework to address this includes methods that assume that the classifier does not have access to the group information, and therefore propose re-weighting high loss examples [29], and data sub-sampling to balance classes and groups [21]. Nevertheless, the methods mentioned above rely on the training and test class sets being identical, making them unsuitable for direct application in zero-shot learning scenarios.
@@ B.1 Derivation of the Loss
We begin by revisiting the parametric model introduced in §3. Let $z_{i}|c_{i}\sim\mathcal{N}\left(c_{i},\Sigma_{z}\right)$, where $\Sigma_{z}=\nu_{z}I_{d}$, $0<\nu_{z}\in\mathbb{R}$, and $I_{d}$ is the $d$ dimensional identity matrix. Classes $c_{i}$ are drawn according to a Gaussian distribution $c_{i}\sim\mathcal{N}\left(0,\Sigma_{a}\right)$ corresponding to their type $a\in\{a_{1},a_{2}\}$. Here, we use a simpler (although less intuitive) notation for the values of the diagonal matrices $\Sigma_{a}$:
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left\|w\left(z_{i}-z_{j}\right)\right\|^{4}=$ $\operatorname{\mathbb{E}}_{y_{ij}=0}\Bigl{[}2\operatorname{tr}\left(\Sigma^{2}\right)+4\operatorname{tr}\left(\Sigma\Sigma_{\mu}\right)+\operatorname{tr}^{2}(\Sigma)+2\operatorname{tr}\left(\Sigma\right)\operatorname{tr}\left(\Sigma_{\mu}\right)\Bigr{.}$
.$\Bigl{.}+2\operatorname{tr}\left(\Sigma_{\mu}^{2}\right)+\left(\operatorname{tr}\left(\Sigma_{\mu}\right)\right)^{2}\Bigr{]}$
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}\left(\Sigma\Sigma_{\mu}\right)\right]=$ $2\nu_{z}\operatorname{tr}\Bigl{(}W^{T}W\left[2\rho^{2}W^{T}\Sigma_{a_{1}}W+2(1-\rho)^{2}W^{T}\Sigma_{a_{2}}W\right.\Bigr{.}$
+.$+\Bigl{.}\left.2\rho(1-\rho)W^{T}\left(\Sigma_{a_{1}}+\Sigma_{a_{2}}\right)W\right]\Bigr{)}$
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}\left(\Sigma_{\mu}\right)\right]=$ $\rho^{2}\operatorname{tr}\left(2W^{T}\Sigma_{a_{1}}W\right)+(1-\rho)^{2}\operatorname{tr}\left(2W^{T}\Sigma_{a_{1}}W\right)$
$+2\rho(1-\rho)\operatorname{tr}\left(W^{T}\left(\Sigma_{a_{1}}+\Sigma_{a_{2}}\right)W\right)$
$=$ $2\left[\nu_{0}\sum_{i=1}^{d_{0}}w_{i}^{2}+\alpha_{1}\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}+\alpha_{2}\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}\right],r$ (27)
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}\left(\Sigma_{\mu}^{2}\right)\right]=$ $\rho^{2}\operatorname{tr}\left((2W^{T}\Sigma_{a_{1}}W)^{2}\right)+(1-\rho)^{2}\operatorname{tr}\left((2W^{T}\Sigma_{a_{2}}W)^{2}\right)$
$+2\rho(1-\rho)\operatorname{tr}\left((W^{T}\left(\Sigma_{a_{1}}+\Sigma_{a_{2}}\right)W)^{2}\right)$
@@
$\operatorname{\mathbb{E}}_{y_{ij}=0}\left[\operatorname{tr}^{2}\left(\Sigma_{\mu}\right)\right]=$ $2\Biggl{[}2\nu_{0}^{2}\left(\sum_{i=1}^{d_{0}}w_{i}^{2}\right)^{2}+\beta_{1}\left(\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}\right)^{2}+\beta_{2}\left(\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}\right)^{2}\Biggr{.}$
$+4\gamma_{0,1}\sum_{i=1}^{d_{0}}w_{i}^{2}\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}+4\gamma_{0,2}\sum_{i=1}^{d_{0}}w_{i}^{2}\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}$
+.$\Biggl{.}4\gamma_{1,2}\sum_{i=d_{0}+1}^{d_{0}+d_{1}}w_{i}^{2}\sum_{i=d_{0}+d_{1}+1}^{d}w_{i}^{2}\Biggr{]},+ .$ (30)
@@
Finally, due to symmetry, at the optimal solution we have
$w_{i}=\begin{cases}u_{0}&0\leq i\leq d_{0}\\ u_{1}&d_{0}+1\leq i\leq d_{0}+d_{1}\\ u_{2}&d_{0}+d_{1}+1\leq i\leq d,\end{cases}$ (32)
@@ B.2 Analysis of the Optimal Solution (Proof of Proposition 1)
Proposition 1 shows that when $d_{1}$ and $d_{2}$ are relatively similar, the optimal solution on the training distribution, assigns more weight to components with high variance in the training data than to those with high variance in the shifted test distribution.
We begin by defining the required condition on $d_{1}$ and $d_{2}$. Denote
@@
$h_{l}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)\coloneqq\frac{\psi_{1}}{\psi_{2}}\frac{\left(\alpha_{2}+\nu_{z}\right)}{\left(\alpha_{1}+\nu_{z}\right)},\quad h_{u}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)\coloneqq\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}},$ (35)
and the corresponding condition
$h_{l}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)<\frac{d_{2}+2}{d_{1}+2}<h_{u}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right).$ (36)
While this condition is sufficient, it is not necessary.Values of $\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}$ and $d_{1},d_{2}$ that satisfy 35 provide an example requiring only a simple analysis, without a full characterization of the optimal solution, for the failure of optimization over the training distribution. However, such failures can occur for additional parameter values, and the full characterization is provided in Appendix B.3.
@@
Denote
$\Delta=$ $\left(2+d_{1}\right)\left[d_{2}u_{2}^{2}\eta_{12}\psi_{1}-\left(\alpha_{2}+\nu_{z}\right)\psi_{1}+d_{0}u_{0}^{2}\psi_{1}\eta_{02}\right]$
-- $\left(2+d_{2}\right)\left[d_{2}u_{2}^{2}\eta_{12}\psi_{2}-\left(\alpha_{1}+\nu_{z}\right)\psi_{2}+d_{0}u_{0}^{2}\psi_{2}\eta_{01}\right],$ (44)
and thus
$d_{1}u_{1}^{2}-d_{2}u_{2}^{2}=$ $\frac{1}{2\left(2+d_{1}\right)\left(2+d_{2}\right)\psi_{1}\psi_{2}}\frac{1}{1-\frac{1}{\xi}\left(2+d_{1}\right)\psi_{1}\eta_{12}}\Delta.$ (45)
Note that for $d_{1},d_{2}$ such that
$\begin{cases}\left(2+d_{1}\right)\psi_{1}-\left(2+d_{2}\right)\psi_{2}>0&\Rightarrow\frac{d_{2}+2}{d_{1}+2}<\frac{\psi_{1}}{\psi_{2}}\\ \left(2+d_{2}\right)\left(\alpha_{1}+\nu_{z}\right)\psi_{2}-\left(2+d_{1}\right)\left(\alpha_{2}+\nu_{z}\right)\psi_{1}>0&\Rightarrow\frac{d_{2}+2}{d_{1}+2}>\frac{\psi_{1}}{\psi_{2}}\frac{\left(\alpha_{2}+\nu_{z}\right)}{\left(\alpha_{1}+\nu_{z}\right)}\\ \left(2+d_{1}\right)\psi_{1}\eta_{02}-\left(2+d_{2}\right)\psi_{2}\eta_{01}>0&\Rightarrow\frac{d_{2}+2}{d_{1}+2}<\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}}\end{cases}$
we have $\Delta>0$. Since $\frac{\eta_{02}}{\eta_{01}}<1$, this reduces to the last two conditions and therefore, in particular for
$\frac{\psi_{1}}{\psi_{2}}\frac{\left(\alpha_{2}+\nu_{z}\right)}{\left(\alpha_{1}+\nu_{z}\right)}<\frac{d_{2}+2}{d_{1}+2}<\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}}$ (46)
we have $\Delta>0$. Additionally, note that
$1-\frac{1}{\xi}\left(2+d_{1}\right)\psi_{1}\eta_{12}$ $=1-\frac{\eta_{12}}{2\left(2+d_{1}\right)\left(2+d_{2}\right)\psi_{1}\psi_{2}}\left(2+d_{1}\right)\psi_{1}=\frac{2\left(2+d_{2}\right)\psi_{2}-\eta_{12}}{2\left(2+d_{2}\right)\psi_{2}}$ (47)
and thus $01-\frac{1}{\xi}\left(2+d_{1}\right)\psi_{1}\eta_{12}>$ 0 iff
$d_{2}+2>\frac{1}{2}\frac{\eta_{12}}{\psi_{2}}.$ (48)
Combining these conditions reduces to
$\frac{\psi_{1}}{\psi_{2}}\frac{\left(\alpha_{2}+\nu_{z}\right)}{\left(\alpha_{1}+\nu_{z}\right)}<\frac{d_{2}+2}{d_{1}+2}<\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}},$ (49)
and therefore, for $\nu_{z},\nu_{0},\nu_{1},\nu_{2},d_{1},d_{2}$ satisfying
$\frac{\psi_{1}}{\psi_{2}}\frac{\left(\alpha_{2}+\nu_{z}\right)}{\left(\alpha_{1}+\nu_{z}\right)}<\frac{d_{2}+2}{d_{1}+2}<\frac{\psi_{1}}{\psi_{2}}\frac{\eta_{02}}{\eta_{01}}.$ (50)
[…]
@@ B.3 Explicit Expression for the Optimal Representation
$\frac{\partial}{\partial\left(w_{i}^{2}\right)}\operatorname{\mathbb{E}}_{y=0}\left[\operatorname{tr}\left(\Sigma\Sigma_{\mu}\right)\right]$ =$=\begin{cases}8\nu_{z}\nu_{0}w_{i}^{2},&1\leq i\leq d_{0}\\ 8\nu_{z}\alpha_{1}w_{i}^{2},&d_{0}+1\leq i\leq d_{0}+d_{1}\\ 8\nu_{z}\alpha_{2}w_{i}^{2},&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (53)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ \frac{\partial\big{[}\operatorname{tr}(\Sigma)\operatorname{\mathbb{E}}_{y=0}[\operatorname{tr}\Sigma_{\mu}]\big{]}}{\partial\left(w_{i}^{2}\right)}=$ (54)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ \begin{cases}4\nu_{z}\left[2\nu_{0}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+\left(\alpha_{1}+\nu_{0}\right)\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+\left(\alpha_{2}+\nu_{0}\right)\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}\right]&1\leq i\leq d_{0}\\ 4\nu_{z}\left[\left(\nu_{0}+\alpha_{1}\right)\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+2\alpha_{1}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+\left(\alpha_{2}+\alpha_{1}\right)\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}\right]&d_{0}+1\leq i\leq d_{0}+d_{1}\\ 4\nu_{z}\left[\left(\nu_{0}+\alpha_{2}\right)\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+\left(\alpha_{1}+\alpha_{2}\right)\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+2\alpha_{2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}\right]&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (55)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ \frac{\partial}{\partial\left(w_{i}^{2}\right)}\operatorname{\mathbb{E}}_{y=0}\left[\operatorname{tr}^{2}\left(\Sigma_{\mu}\right)\right]=$ (57)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ \begin{cases}8\nu_{0}^{2}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+8\gamma_{0,1}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+8\gamma_{0,2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}&1\leq i\leq d_{0}\\ 8\gamma_{0,1}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+4\beta_{1}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+8\gamma_{1,2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}&d_{0}+1\leq i\leq d_{0}+d_{1}\\ 8\gamma_{0,2}\sum\limits_{j=1}^{d_{0}}w_{j}^{2}+8\gamma_{1,2}\sum\limits_{j=d_{0}+1}^{d_{0}+d_{1}}w_{j}^{2}+4\beta_{2}\sum\limits_{j=d_{0}+d_{1}+1}^{d}w_{j}^{2}&d_{0}+d_{1}+1\leq i\leq d\end{cases}$ (58)
@@
Thus, we can write for the symmetric solution
$\partial_{0}$ $=-2m\left(\nu_{0}+\nu_{z}\right)+u_{0}^{2}G_{0,0}+u_{1}^{2}G_{0,1}+u_{2}^{2}G_{0,2},$ (59)
$\partial_{1}$ $=-2m\left(\alpha_{1}+\nu_{z}\right)+u_{0}^{2}G_{1,0}+u_{1}^{2}G_{1,1}+u_{2}^{2}G_{1,2},$ (60)
$\partial_{2}$ $=-2m\left(\alpha_{2}+\nu_{z}\right)+u_{0}^{2}G_{2,0}+u_{1}^{2}G_{2,1}+u_{2}^{2}G_{2,2},$ (61)